((xn,yn) is convergent iff both (xn}) and (yn) are convergent. In fact, for(x0 ,y0) in R2, we have (xn ,yn) converging to (x0,y0) iff xn converges to x0 and yn converges to y0

1
Expert's answer
2022-04-19T03:35:23-0400

ANSWER

To prove the statement, we use the definitions:

Definition 1 Let "\left\{ \left( { x }_{ n\ },{y}_{n} \right) \right\}" is a sequence in "\R^{2}" . We say that "\left( { x }_{ n\quad },{ y }_{ n } \right)" convergens to "\left( { x }_{0\ },{ y }_{ 0 } \right)" and write "\left( { x }_{ n\ },{ y }_{ n } \right) \rightarrow \left( { x }_{ 0\ },{ y }_{ 0 } \right)" if for every "\varepsilon >0\quad"there is an "N_{0}\in \N" such that for all "n\in \N," if "n>N_{0}" , then

"\left\| \left( { x }_{ n\quad },{ y }_{ n } \right) -\left( { x }_{ 0\quad },{ y }_{ 0 } \right) \right\| =\sqrt { { \left( { x }_{ n }-{ x }_{ 0\quad } \right) }^{ 2 }+{ \left( { y }_{ n }-{ y }_{ 0\quad } \right) }^{ 2 } } <\varepsilon"

Definition 2 Let "\left\{ \left( { x }_{ n }\right ) \right\}" is a sequence in "\R" . We say that "\left( { x }_{ n } \right)" convergens to "\left( { x }_{0\ } \right)" and write "{ x }_{ n } \rightarrow { x }_{ 0 }" if for every "\varepsilon >0\quad"there is an "N_{x}\in \N" such that for all "n\in\N" if "n>N_{x}" , then

"|x_{ n }-x_{ 0 }|<\varepsilon"

1) Let "\left( { x }_{ n\ },{ y }_{ n } \right) \rightarrow \left( { x }_{ 0\ },{ y }_{ 0 } \right)" .If "\varepsilon>0" and "n>N_{0}" , such that "\left\| \left( { x }_{ n\quad },{ y }_{ n } \right) -\left( { x }_{ 0\quad },{ y }_{ 0 } \right) \right\| <\varepsilon" ,then for all "n>N_{0}" :

"|{ x }_{ n }-{ x }_{ 0\quad }|\quad =\sqrt { { \left( { x }_{ n }-{ x }_{ 0\quad } \right) }^{ 2 }\quad } \le \sqrt { { \left( { x }_{ n }-{ x }_{ 0\quad } \right) }^{ 2 }+{ \left( { y }_{ n }-{ y }_{ 0\quad } \right) }^{ 2 } } =\left\| \left( { x }_{ n\quad },{ y }_{ n } \right) -\left( { x }_{ 0\quad },{ y }_{ 0 } \right) \right\|<\varepsilon"

and "|{ y }_{ n }-{ y }_{ 0\ }|\quad =\sqrt { { \left( { y }_{ n }-{ y }_{ 0\quad } \right) }^{ 2 } } \le \sqrt { { \left( { x }_{ n }-{ x }_{ 0\quad } \right) }^{ 2 }+{ \left( { y }_{ n }-{ y }_{ 0\quad } \right) }^{ 2 } } =\left\| \left( { x }_{ n\quad },{ y }_{ n } \right) -\left( { x }_{ 0\quad },{ y }_{ 0 } \right) \right\| <\varepsilon"

Therefore, by the definition 2 "{ x }_{ n } \rightarrow { x }_{ 0 }" and "{y }_{ n } \rightarrow { y }_{ 0 }"

2) If "{ x }_{ n } \rightarrow { x }_{ 0 }" and "{y }_{ n } \rightarrow { y }_{ 0 }" , then by the definition 2 , for every "\varepsilon >0\quad"there is an "N_{x}, N_{y}\in \N" such that for all "n\in \N," if "n>N_{x}" , then

"|x_{n} -x_{0}|<\frac { \varepsilon }{ \sqrt { 2 } }"

and if "n>N_{y}" , then

"|y_{n} -y_{0}|<\frac { \varepsilon }{ \sqrt { 2 } }" .

Let "{ N }_{ 0 }=\max { \left\{ N_{ x },\quad N_{ y } \right\} }" . Since "N_{0}>N_{x}, N_{0}>N_{y}" , then for all "n>N_{0}" "(\Rightarrow n>N_{x}, n>N_{y})"

"\left\| \left( { x }_{ n\quad },{ y }_{ n } \right) -\left( { x }_{ 0\quad },{ y }_{ 0 } \right) \right\| =\sqrt { { \left( { x }_{ n }-{ x }_{ 0\quad } \right) }^{ 2 }+{ \left( { y }_{ n }-{ y }_{ 0\quad } \right) }^{ 2 } } <\sqrt { \frac { { \varepsilon }^{ 2 } }{ 2 } +\frac { { \varepsilon }^{ 2 } }{ 2 } } =\varepsilon" .

Therefore, by the definition 1, "\left( { x }_{ n\ },{ y }_{ n } \right) \rightarrow \left( { x }_{ 0\ },{ y }_{ 0 } \right)" .

Hence, "1)\Leftrightarrow 2)" .